A POT OF WATER HEATED BY TRANSFERRING 1676 KJ OF HEAT ENERGY TO THE WATER IF THERE 0D 0AIS 5.0 KG OF WATER IN THE POT , THE TEMPARATURE IS RAISED BY 80.0 K , WHAT IS THE SPECIFIC HEAT OF WATER


1
Expert's answer
2021-09-04T18:21:22-0400
"Q=mc\Delta T,""c=\dfrac{Q}{m\Delta T}=\dfrac{1676\cdot10^3\ J}{5\ kg\cdot80\ K}=4190\ \dfrac{J}{kg\cdot \!K}"